# LongCaptioning: Unlocking the Power of Long Caption Generation in Large Multimodal Models

### v1

@@ Abstract
Large multimodal models (LMMs) have shown remarkable performance in video understanding tasks and can even process videos longer than one hour. However, despite their ability to handle long inputs, generating outputs with corresponding levels of richness remains a challenge. In this paper, we explore the issue of long outputs in LMMs using video captioning as a proxy task, and we find that open-source LMMs struggle to consistently generate outputs exceeding about 300 words. Through controlled experiments, we find that the scarcity of paired examples with long-captions during training is the primary factor limiting the model’s output length. However, manually annotating long-caption examples is time-consuming and expensive. To address this, we propose the LongCaption-Agent, a framework that synthesizes long caption data by aggregating multi-level descriptions. Using LongCaption-Agent, we curated a new long-caption dataset, LongCaption-10K. We also develop LongCaption-Bench, a benchmark designed to comprehensively evaluate the quality of long captions generated by LMMs. By incorporating LongCaption-10K into training, we enable LMMs to generate captions exceeding 1,000 words, while maintaining high output quality. In LongCaption-Bench, our 8B parameter model achieved state-of-the-art performance, even surpassing larger proprietary models. We will release the dataset and code after publication.
@@ 1 Introduction
Large multimodal models (LMMs) [47, 42, 28, 4, 25, 18, 33, 41, 26] have made significant progress in video understanding tasks. Given the rich information in videos, recent works have attempted to handle longer visual context inputs. For example, LongVA [47] extended the input video length to 1 hour, which corresponds to approximately 10,000 visual tokens. However, despite their ability to handle long input, they still face challenges in generating long captions. Here, we distinguish between dense video captioning [21, 43, 19, 38] and long caption generation. The former aims to identify different event timestamps within a video and generate corresponding captions for each event. This task still focuses on short video and short caption scenarios, with the generated captions typically not exceeding 30 words. In contrast, long caption generation requires producing captions that exceed or approach 1,000 words.
In this paper, we explore the issue of long caption generation in LMMs. As shown in Figure 2, we collect 5 movie videos, each approximately 20 minutes, and segment each video into 6 clips of 1, 2, 5, 10, 15, and 20 minutes. For videos of different duration, we sample one frame every 4 seconds as an input to the model. This ensures that as the video duration increases, the amount of information it contains also becomes richer. Then, we use the same prompts (detailed prompts are provided in the Appendix) to have different models perform the video captioning task. We find that although the amount of information in the videos increases, the open-source LMMs consistently fail to generate outputs exceeding about 300 words, leading to captions that lose most of the visual information. In contrast, the proprietary model ($i.e.$, Gemini 1.5 Pro [34]) performs exceptionally well, easily generating descriptions of over 1,000 words for 20-minute videos, thereby preserving rich fine-grained information. However, considering the lack of open-source technical solutions for proprietary models and the challenges they face in deployment on edge devices, we aim to explore a low-cost method to enhance the long caption generation capabilities of open-source LMMs.
@@
Figure 2: Visualization of model output length when trained on datasets (each containing 5,000 samples) with varying caption lengths and sampled frame counts. All captions are synthetically generated.
LMMs typically consist of a visual encoder, a modality projector, and a language decoder, where the language decoder is usually a pretrained large language model (LLM). As a core component, the development of LLMs often drives advancements in LMMs. LongWriter [5] is the first to explore the long text generation problem in LLMs. By constructing a supervised fine-tuning dataset for long texts, it successfully extended the model output from 2,000 to 10,000 words. In this paper, we use the video captioning task as a proxy to explore long-text output in LMMs for the first time. We conduct controlled experiments to investigate the key factors that affect the output length of LMMs.
[…]
Figure 3: The LongCaption-Agent framework. The framework uses an off-the-shelf LMM ($i.e.$, MiniCPMV2.6 8B [45]) to first generate frame-level descriptions for each sampled frame, and then iteratively produce clip-level descriptions for each clip. Finally, an off-the-shelf LLM ($i.e.$, GLM4-Long [12]) is used to integrate the frame-level and clip-level descriptions into a complete video long-caption.
To address this, we propose the LongCaption-Agent, an automated framework for synthesizing long-caption data. This framework utilizes off-the-shelf LMMs and LLMs to synthesize captions in three stages: frame-level, clip-level, and video-level. By employing multi-level information extraction and summarization, the framework produces comprehensive long-captions. Building upon the LongCaption-Agent, we develop the LongCaption-10K dataset, which includes 10,000 long-caption examples. By incorporating the LongCaption-10K dataset into training, we extend the output length of LMM to over 1,000 words, successfully unlocking its long-caption generation capabilities. Additionally, to further enable the model to handle longer sampled frame inputs during inference, we introduce a visual context window extension method [39] in the training phase. This method effectively mitigates the issue of reduced output length when transitioning from short-sequence training to long-sequence inference, as demonstrated in Figure 2.
To rigorously evaluate the long-caption generation capabilities of LMM, we develop LongCaption-Bench, a comprehensive benchmark for assessing the quality of long-captions generated by LMMs. It includes 281 test videos with an average duration of 1060.4 seconds. Each video has a global description generated by Gemini 1.5 Pro [34], which is then manually reviewed and modified. The average caption length is 1161.3 words.
@@ 2 Related Work
Large Multimodal Model Recent advancements in large language models (LLMs) [35, 17, 44, 3, 32, 1] have demonstrated impressive language understanding and generation capabilities. This success has sparked interest in large multimodal models (LMMs) [23, 10, 46, 45], which typically consist of visual encoders, modality projectors, and pretrained language model decoders. LMMs initially made breakthroughs in image understanding tasks. With the construction of high-quality video-text datasets, more researchers are applying LMMs to video understanding tasks [47, 42, 28, 23, 25, 18, 33, 41, 26]. For example, models like VideoChatGPT [29], VideoChat [24], and Video-LLaMA [9] have enhanced the video understanding capabilities of LMMs through high-quality data and fine-tuning techniques. These models have shown excellent performance in short video understanding. Recently, some efforts [47, 39, 42, 28] have been made to input long videos into LMMs, achieving some progress. For instance, MovieChat [33] introduced a memory mechanism to compress long video tokens into a fixed size. Additionally, LongVA [47] extended the context window by continuously training LLMs on long texts. Although they perform well in long video-QA tasks, they face challenges in generating video captions that require global descriptions.
Video Captioning Early video captioning methods used template-based approaches [20, 13, 22], which lacked flexibility. With the development of deep learning, expert models based on CNN-RNN and Transformer architectures replaced previous methods [11, 15]. However, these approaches typically handle only short videos of a few seconds, and the generated captions are also brief. The VideoReCap [16] model was the first to attempt generating descriptions for long videos recursively. However, its descriptions often do not exceed 100 words for a 60-minute video, inevitably missing much of the video content. Recently, some efforts [8, 7] have attempted to combine LMMs and LLMs to construct large-scale video captioning datasets. For instance, Panda70M [8] constructed semantically consistent videos by splitting and merging based on semantic understanding, and then used a pre-trained model to generate captions for each video. ShareGPT4Video [7] proposed a differential video captioning strategy, leveraging GPT-4V [1] to synthesize video captions by identifying differences between adjacent frames. However, these methods typically focus on short video-caption examples.
@@ 3 LongCaption-Agent: A Long-Caption Synthesis Framework
Table 1: Existing Video-Text Datasets
@@ 3.1 Framework
In Section 1, we analyze the challenges faced by large multimodal models (LMMs) in generating long captions, identifying the scarcity of long-caption examples during training as a key factor. A statistical review of commonly used video-text datasets, shown in Table 1, reveals that their average text lengths are all under 100 words. For instance, Pand70M [8] synthesizes video captions using off-the-shelf multimodal models, combining outputs from various teacher models and a retrieval model to select the best match. However, due to output length limitations, the generated captions remain relatively short. Additionally, synthesizing long captions with proprietary LMMs or manual annotation is prohibitively expensive.
@@
Step 1 - Frame-Level Captioning: Considering that current open-source LMMs struggle to generate long-captions for videos, we instead sample frames as model inputs to extract fine-grained static information from each frame. As shown in Figure 3, in this work, we use MiniCPMV2.6 8B [45] as the frame-level caption generation model. Some studies have attempted to use LLMs to summarize frame-level captions into video captions. However, by neglecting the temporal information of the original video, this approach often fails to accurately capture the dynamic changes in events and the continuity of actions, resulting in captions that lack coherence and completeness. Temporal information is crucial for understanding causal relationships, action sequences, and scene transitions in videos. Relying solely on frame-level summaries tends to produce fragmented captions that struggle to reflect the overall semantics of the video. Therefore, in this work, we use frame-level captions only as a supplementary source of fine-grained static information for clip-level captions.
@@
$\operatorname{caption}_{t}=\operatorname{LMM}(\operatorname{clip}_{t},\quad\operatorname{caption}_{t-1})$ (1)
@@
Specifically, our pipeline is divided into the following steps: 1) Timestamp Annotation: First, we annotate each clip-level caption and each frame-level caption with a timestamp, allowing the LLM to align the frame-level descriptions with the clip-level captions in terms of time. 2) Frame-Level Caption Grouping: Next, based on the timestamps of clip-level captions, we group the continuous frame-level captions into different sets, with each set corresponding to the time range of a clip. This ensures that the frame-level captions can be summarized within a local time range without losing details. 3) Interleaved Input Combination: Finally, we interleave all the frame-level caption sets with the clip-level captions and, combined with carefully designed prompts, feed them as input to the LLM. The prompts are designed to guide the LLM in effectively merging the fine-grained information from the frame-level captions with the temporal information from the clip-level captions, thereby generating a coherent and information-rich long-caption. The formula can be expressed as:
@@
3.2 LongCaption-10K Dataset

### v2

# LongCaptioning: Unlocking the Power of Long Video Caption Generation in Large Multimodal Models

@@ Abstract
Large Multimodal Models (LMMs) have demonstrated exceptional performance in video captioning tasks, particularly for short videos. However, as the length of the video increases, generating long, detailed captions becomes a significant challenge. In this paper, we investigate the limitations of LMMs in generating long captions for long videos. Our analysis reveals that open-source LMMs struggle to consistently produce outputs exceeding 300 words, leading to incomplete or overly concise descriptions of the visual content. This limitation hinders the ability of LMMs to provide comprehensive and detailed captions for long videos, ultimately missing important visual information. Through controlled experiments, we find that the scarcity of paired examples with long-captions during training is the primary factor limiting the model’s output length. However, manually annotating long-caption examples for long-form videos is time-consuming and expensive. To overcome the annotation bottleneck, we propose the LongCaption-Agent, a framework that synthesizes long caption data by hierarchical semantic aggregation. Using LongCaption-Agent, we curated a new long-caption dataset, LongCaption-10K. We also develop LongCaption-Bench, a benchmark designed to comprehensively evaluate the quality of long captions generated by LMMs. By incorporating LongCaption-10K into training, we enable LMMs to generate captions exceeding 1,000 words for long-form videos, while maintaining high output quality. In LongCaption-Bench, our model achieved State-of-The-Art performance, even surpassing larger proprietary models like GPT4o.
@@ 1 INTRODUCTION
Corresponding author: Zhenzhong Chen, E-mail:zzchen@ieee.org
Video captioning, a critical task in multimodal understanding, requires models to generate comprehensive and context-rich descriptions for video content. With the rapid development of Large Multimodal Models (LMMs) [1, 2, 3, 4, 5, 6, 7, 8, 9], video captioning has achieved significant progress. It generates descriptions for videos by concatenating visual features and textual features as input to large language models. However, previous research has predominantly focused on short videos, where the scope of context is limited, and captions typically remain concise. In contrast, long videos introduce additional challenges: they contain more complex, diverse content that requires models to capture and maintain long-term dependencies across frames and segments. This increased temporal span makes it more difficult for models to generate captions that are not only comprehensive but also coherent and contextually consistent over extended periods.
To evaluate the performance of current LMMs on long video captioning, we design a set of experiments specifically tailored to this task. In particular, we collect 5 movie videos, each approximately 20 minutes long, and segment each video into 6 clips of 1, 2, 5, 10, 15, and 20 minutes. For videos of different durations, we sample one frame every 4 seconds as input to the model. This ensures that as the video duration increases, the amount of information it contains also becomes richer. We then use the same prompts (detailed prompts are provided in the Appendix) to have different models perform the video captioning task. We find that although the amount of visual information in the videos increases with length, open-source LMMs consistently struggle to generate outputs exceeding approximately 300 words, even as video duration and complexity grow. This limitation results in a loss of critical visual details, as the models fail to capture or adequately describe the full scope of content. For example, as shown in Figure 1 (a), Qwen-VL 72B generated only 79 words for a 17-minute video, significantly omitting visual details. Figure 1 (b) shows that other open-source models exhibit a similar pattern, with outputs consistently below 300 words even as video length increases. In contrast, the proprietary model (e.g., Gemini 1.5 Pro [10]) performs exceptionally well, generating over 1,000 words for a 20-minute video, effectively preserving rich, fine-grained information. However, given the lack of open-source solutions for such proprietary models and the deployment challenges they pose on edge devices, we aim to explore a low-cost method to enhance the long-form video caption generation capabilities of open-source LMMs.
@@
LMMs typically consist of a visual encoder, a modality projector, and a language decoder, where the language decoder is usually a pretrained large language model (LLM). As a core component, the development of LLMs often drives advancements in LMMs. LongWriter [11] is the first to explore the long text generation problem in LLMs. By constructing a supervised fine-tuning dataset for long texts, it successfully extended the model output from 2,000 to 10,000 words. Inspired by this work, we design a set of controlled experiments to investigate long caption generation in the context of long video captioning. We find that despite the increasing volume of visual information with longer videos, open-source LMMs consistently fail to generate captions exceeding 300 words. The root cause of this limitation lies in the scarcity of long-caption examples during training, which hinders the model’s ability to generate outputs beyond a certain length. This issue becomes particularly pronounced as the video duration increases, as the model’s context window is insufficient to capture all relevant details. To overcome this limitation, one potential solution is to construct datasets of video-long caption pairs. However, manually annotating such datasets is time-consuming and expensive, which presents a significant barrier to improving long-caption generation in LMMs.
To address this, we propose the LongCaption-Agent, an automated framework for synthesizing long-caption data. This framework utilizes off-the-shelf LMMs and LLMs to synthesize captions in three stages: frame-level, clip-level, and video-level. By employing multi-level information extraction and summarization, the framework produces comprehensive long-captions. Building upon the LongCaption-Agent, we develop the LongCaption-10K dataset, which includes 10,000 long-caption examples. By incorporating the LongCaption-10K dataset into training, we extend the output length of LMM to over 1,000 words, successfully unlocking its long video caption generation capabilities. Additionally, to further enable the model to handle longer sampled frame inputs during inference, we introduce a visual context window extension method [12] in the training phase. This method effectively mitigates the issue of reduced output length when transitioning from short-sequence training to long-sequence inference.
Traditional $n$-gram-based metrics ($e.g.$, CIDEr [13]) commonly used for image and video captioning fall short for long captions due to the flexibility of language. To rigorously evaluate the long-caption generation capabilities of LMM, we develop LongCaption-Bench, a comprehensive benchmark for assessing the quality of long-captions generated by LMMs. It includes 281 test videos with an average duration of 1060.4 seconds. Each video has a global description generated by Gemini 1.5 Pro [10], which is then manually reviewed and modified. The average caption length is 1161.3 words.
@@
The rest of the paper is organized as follows. In Section 2, we provide an overview of related work. Section 3 presents a detailed description of the data synthesis process for the LongCaption-Agent and the statistical characteristics of LongCaption-10K dataset. In Section 4, we introduce LongCaption-Bench, the evaluation benchmark for long video captioning. Section 5 elaborates on the experimental setup and training details of LongCaptioning-8B, followed by an analysis of the experimental results. Finally, Section 6 provides a summary of the paper.
Figure 2: The LongCaption-Agent framework. The framework uses an off-the-shelf LMM ($i.e.$, MiniCPMV2.6-8B [14]) to first generate frame-level descriptions for each sampled frame, and then iteratively produce clip-level descriptions for each clip. Finally, an off-the-shelf LLM ($i.e.$, GLM4-Long [15]) is used to integrate the frame-level and clip-level descriptions into a complete video long-caption, incorporating additional context such as timestamps.
@@ 2 Related Work
2.1 Large Multimodal Model
Recent advancements in large language models (LLMs) [16, 17, 18, 19, 20, 21] have demonstrated impressive language understanding and generation capabilities. This success has sparked interest in large multimodal models (LMMs) [22, 23, 24, 14], which typically consist of visual encoders, modality projectors, and pretrained language model decoders. LMMs initially made breakthroughs in image understanding tasks. With the construction of high-quality video-text datasets, more researchers are applying LMMs to video understanding tasks [1, 2, 3, 22, 5, 6, 7, 8, 9]. For example, models like VideoChatGPT [25], VideoChat [26], and Video-LLaMA [27] have enhanced the video understanding capabilities of LMMs through high-quality data and fine-tuning techniques. These models have shown excellent performance in short video understanding. Recently, some efforts [1, 12, 2, 3] have been made to input long videos into LMMs, achieving some progress. For instance, MovieChat [7] introduced a memory mechanism to compress long video tokens into a fixed size. Additionally, LongVA [1] extended the context window by continuously training LLMs on long texts. Although they perform well in long video-QA tasks, they face challenges in generating video captions that require global descriptions. For a 20-minute video, they struggle to output even 300 words of description. This limits the application of the model in video understanding.
2.2 Video Captioning
Early video captioning methods used template-based approaches [28, 29, 30], which lacked flexibility. With the development of deep learning, expert models based on CNN-RNN and Transformer architectures replaced previous methods [31, 32]. However, these approaches typically handle only short videos of a few seconds, and the generated captions are also brief. The VideoReCap [33] model was the first to attempt generating descriptions for long videos recursively. However, its descriptions often do not exceed 100 words for a 60-minute video, inevitably missing much of the video content. Dense video captioning [34, 35, 36] typically identifies different event timestamps within a video and generates corresponding captions for each event. However, these methods still focus on short video-short caption scenarios, with annotated captions generally limited to no more than 30 words. Recently, some efforts [37, 38] have attempted to combine LMMs and LLMs to construct large-scale video captioning datasets. For instance, Panda70M [37] constructed semantically consistent videos by splitting and merging based on semantic understanding, and then used a pre-trained model to generate captions for each video. ShareGPT4Video [38] proposed a differential video captioning strategy, leveraging GPT-4V [21] to synthesize video captions by identifying differences between adjacent frames. However, these methods typically focus on short video-caption examples.
@@ 3 LongCaption-Agent: A Long-Caption Synthesis Framework
Figure 3: Existing Video-Text Datasets
@@ 3.1 Framework
In Section 1, we conduct a detailed analysis of the challenges faced by previous large multimodal models (LMMs) in generating long-captions. Our investigation reveals that the primary factor contributing to these challenges is the scarcity of long-caption examples during the training phase. We conduct a statistical analysis of the commonly used video-text datasets. Table 4 shows the average video duration and average text length of these datasets, with all of their average text lengths being less than 100. Among them, Pand70M [37] uses off-the-shelf open-source multimodal models to synthesize video caption data. Specifically, it employs multiple multimodal teacher models to generate captions for different modality combinations, and then uses a trained retrieval model to select the caption that best matches the video. However, due to the limitation of the model’s output length, the captions synthesized by Pand70M are relatively short. In addition, using proprietary large multimodal models (e.g., Gemini 1.5 Pro [10]) to synthesize long-captions, as well as manually annotating long-captions, are both extremely costly.
@@
Step 1 - Frame-Level Captioning: Considering that current open-source LMMs struggle to generate long-captions for videos, we instead sample frames as model inputs to extract fine-grained static information from each frame. As shown in Figure 2, in this work, we use MiniCPMV2.6-8B [14] as the frame-level caption generation model. Some studies have attempted to use LLMs to summarize frame-level captions into video captions. However, by neglecting the temporal information of the original video, this approach often fails to accurately capture the dynamic changes in events and the continuity of actions, resulting in captions that lack coherence and completeness. Temporal information is crucial for understanding causal relationships, action sequences, and scene transitions in videos. Relying solely on frame-level summaries tends to produce fragmented captions that struggle to reflect the overall semantics of the video. Therefore, in this work, we use frame-level captions only as a supplementary source of fine-grained static information for clip-level captions.
@@
$\operatorname{caption}_{t}=\operatorname{LMM}(\operatorname{clip}_{t},\operatorname{caption}_{t-1})$ (1)
@@
Figure 5: The framework of the Large Multimodal Model (LMM). The model consists of a vision encoder that processes video frames and generates vision tokens. These tokens are then passed through a compression layer that reduces the number of vision tokens. The compression layer uses weighted sum and attention mechanisms, incorporating vision tokens, and learnable queries tokens, where the number of learnable queries is much smaller than the number of value and key. The processed information is fed into a Large Language Model (LLM), which leverages LoRA (Low-Rank Adaptation) to train language model.
@@
Specifically, our pipeline is divided into the following steps: 1) Timestamp Annotation: First, we annotate each clip-level caption and each frame-level caption with a timestamp, allowing the LLM to align the frame-level descriptions with the clip-level captions in terms of time. 2) Frame-Level Caption Grouping: Next, based on the timestamps of clip-level captions, we group the continuous frame-level captions into different sets, with each set corresponding to the time range of a clip. This ensures that the frame-level captions can be summarized within a local time range without losing details. 3) Interleaved Input Combination: Finally, we interleave all the frame-level caption sets with the clip-level captions and, combined with carefully designed prompts, feed them as input to the LLM. The prompts are designed to guide the LLM in effectively merging the fine-grained information from the frame-level captions with the temporal information from the clip-level captions, thereby generating a coherent and information-rich long-caption.
@@ 3.2 LongCaption-10K Dataset
Table 2: Evaluatio results of the video-caption relevance score on the LongCaption-Bench across different duration intervals, where “Overall” refers to the average score for the entire dataset, “[300s, 600s)” refers to the score for videos with duration between 300 seconds and 600 seconds, “[600s, 900s)” refers to the score for videos with duration between 600 seconds and 900 seconds, and “[900s, 1200s)” refers to the score for videos with duration between 900 seconds and 1200 seconds, and “[1200s, 1800s)” refers to the score for videos with duration between 1200 seconds and 1800 seconds. The score is calculated by averaging the scores of each video in the corresponding duration interval.
Overall [300s, 600s) [600s, 900s) [900s, 1200s) [1200s, 1800s)
Proprietary models
GPT-4o [20] 2.50 2.78 2.41 2.57 2.50
Open-source models
PLLaVA 7B [8] 1.10 1.09 1.07 1.10 1.18
PLLaVA 32B [8] 1.18 1.21 1.10 1.20 1.17
LongVA 7B [1] 1.15 1.11 1.07 1.17 1.18
LLaVA-Video 7B [46] 1.30 1.33 1.21 1.38 1.24
MiniCPMV2.6-8B [14] 2.18 2.11 2.02 2.24 2.20
Qwen2-VL 7B [47] 1.36 1.22 1.12 1.47 1.36
Qwen2-VL 72B [47] 1.71 1.72 1.66 1.72 1.68
Our trained models
LongCaptioning-8B 2.59 2.56 2.53 2.62 2.58
Table 3: Evaluation results of the length score $S_{l}$ and the quality score $S_{q}$ on the LongCaption-Bench across different duration intervals. Here, $S_{l}$ represents the score for caption length, $S_{q}$ represents the quality score.
Overall [300s, 600s) [600s, 900s) [900s, 1200s) [1200s, 1800s)
$S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$
Proprietary models
GPT-4o [20] 33.6 79.4 64.9 81.5 41.2 81.2 30.2 79.4 24.3 78.7
Open-source models
PLLaVA 7B [8] 4.5 63.2 4.0 64.1 2.3 63.7 3.7 64.9 6.8 62.8
PLLaVA 32B [8] 5.3 65.1 4.9 66.4 2.6 64.7 4.9 65.2 7.7 64.0
LongVA 7B [1] 2.1 64.5 0.0 65.2 0.0 65.5 1.9 64.6 4.0 63.6
LLaVA-Video 7B [46] 4.5 66.8 8.6 67.8 1.8 69.0 3.1 67.2 8.4 64.8
MiniCPMV2.6-8B [14] 11.8 73.5 27.6 72.6 12.8 74.9 10.4 73.1 11.5 73.0
Qwen2-VL 7B [47] 7.7 57.8 0.0 58.9 6.6 56.4 6.4 58.5 11.2 57.6
Qwen2-VL 72B [47] 9.4 70.2 11.8 69.1 7.7 68.6 7.0 71.4 10.1 70.2
Our trained models
LongCaptioning-8B 40.9 81.2 88.2 80.7 53.6 81.7 38.8 81.6 30.3 80.4
Table 4: Ablation study results of the video-caption relevance score on the LongCaption-Bench. ‘w/o Visual Context Window Ext.’ represents the LongCaptioning 8B model trained without the visual context window extension, ‘w/o Clip-Level Captioning’ indicates that only frame-level captions were used during long caption synthesis, ‘w/o Frame-Level Captioning’ refers to the use of only clip-level captions during long caption synthesis, and ‘w/o LongCaption-10k data’ refers to the results without using the LongCaption-10K dataset for training.
Overall [300s, 600s) [600s, 900s) [900s, 1200s) [1200s, 1800s)
LongCaptioning-8B 2.59 2.56 2.53 2.62 2.58
   w/o Visual Context Window Ext. 2.35 2.44 2.13 2.37 2.45
   w/o Clip-Level Captioning 2.21 2.25 2.06 2.28 2.30
   w/o Frame-Level Captioning 2.20 2.20 2.06 2.27 2.28
   w/o LongCaption-10k data 2.18 2.11 2.02 2.24 2.20
Table 5: Ablation study results of the length score $S_{l}$ and quality score $S_{q}$ on the LongCaption-Bench.
Overall [300s, 600s) [600s, 900s) [900s, 1200s) [1200s, 1800s)
$S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$
LongCaptioning-8B 40.9 81.2 88.2 80.7 53.6 81.7 38.8 81.6 30.3 80.4
   w/o Visual Context Window Ext. 32.3 79.1 90.3 76.3 37.5 78.7 30.1 79.7 25.8 78.6
   w/o Clip-Level Captioning 28.7 75.0 64.3 74.8 25.8 76.5 21.0 76.7 19.1 74.9
   w/o Frame-Level Captioning 25.7 74.2 58.8 74.6 21.5 76.1 18.9 76.0 16.0 74.3
   w/o LongCaption-10k data 11.8 73.5 27.6 72.6 12.8 74.9 10.4 73.1 11.5 73.0
Table 6: Performance evaluation on VideoMME [48] benchmark, where “Short” refers to the duration of 0s to 120s, “Medium” refers to the duration of 240s to 900s, “Long” refers to the duration of 1800s to 3600s, and “Overall” refers to the average of all duration ranges.
Methods Short Medium Long Overall
ST-LLM-7B [49] 45.7 36.8 31.3 37.9
VideoLLaVA-7B [9] 45.3 38.0 36.2 39.9
VideoChat2-Mistral-7B [50] 48.3 37.0 33.2 39.5
Chat-UniVi-V1.5-7B [51] 45.7 40.3 35.8 40.6
Qwen-VL-Chat-7B [52] 46.9 38.7 37.8 41.1
VideoLLaMA2-7B [27] 56.0 45.4 42.1 47.9
LLaVA-NeXT-Qwen2-7B [46] 58.0 47.0 43.4 49.5
LongVA-7B [1] 61.1 50.4 46.2 52.6
LLaVA-OneVision-7B [22] 69.3 55.1 49.7 58.2
MiniCPMV2.6-8B [14] 71.3 59.4 51.8 60.9
LongCaptioning-8B 72.2 59.9 53.4 62.4